# The uniqueness theorem for Kasparov theory

### v1

@@ Abstract.
Answering a question of Carrión et al in their recent landmark paper on $\mathrm{C}^{*}$-algebra classification, we prove a general uniqueness theorem for $KK$-theory. Given arbitrary separable $\mathrm{C}^{*}$-algebras $A$ and $B$ and a Cuntz pair consisting of two absorbing representations $\varphi,\psi:A\to\mathcal{M}(B\otimes\mathcal{K})$, the induced element of $KK(A,B)$ vanishes if and only if $\varphi$ and $\psi$ are strongly asymptotically unitarily equivalent. This improves upon the Lin–Dadarlat–Eilers stable uniqueness theorem. The conclusion is deduced by first showing the $K_{1}$-injectivity of an auxiliary $\mathrm{C}^{*}$-algebra associated to the $\mathrm{C}^{*}$-pair $(A,B)$, which is sometimes called the Paschke dual algebra in the literature. Most of the article is concerned with the treatment of an umbrella theorem, which yields such a uniqueness theorem for other variants of $KK$-theory. This encompasses nuclear $KK$-theory, ideal-related $KK$-theory, equivariant $KK$-theory, or any combinations thereof.
@@ Introduction
Kasparov theory, also known as $KK$-theory, is a very powerful invariant for $\mathrm{C}^{*}$-algebras. It is a bivariant version of topological $K$-theory, which includes ordinary $K$-theory and $K$-homology, alongside with Brown–Douglas–Fillmore theory [5], as special cases. Invented by Kasparov [30, 31], it can be viewed as either a bifunctor from (separable) C*-algebras to abelian groups, or as a way to equip the class of (separable) $\mathrm{C}^{*}$-algebras with a more flexible notion of homomorphisms. In fact, given a second-countable locally compact group $G$, Kasparov’s construction also yields such a bifunctor on the category of all (separable) $G$-$\mathrm{C}^{*}$-algebras, which provided impactful applications on classical problems such as the Novikov conjecture. The original approach by Kasparov, commonly referred to as the Kasparov picture, is still the most common way of treating this subject, but there are other equivalent appproaches that are geared towards different types of applications.
In this article we are mainly interested in the Cuntz picture [8] and the Cuntz–Thomsen picture [42] for (equivariant) $KK$-groups. In Cuntz’ framework, elements in Kasparov’s group $KK(A,B)$ for separable $\mathrm{C}^{*}$-algebras $A$ and $B$ can be described as homotopy classes of so-called Cuntz pairs, i.e., pairs of *-homomorphisms $\varphi,\psi:A\to\mathcal{M}(B\otimes\mathcal{K})$ such that $\varphi(a)-\psi(a)\in B\otimes\mathcal{K}$ for all $a\in A$. For equivariant $KK$-theory, one considers pairs of cocycle representations satisfying a similar condition; see also [41, Section 6]. A groundbreaking theorem concerning the analytical properties of $KK$-theory is the so-called stable uniqueness theorem proved independently by Lin [34] and Dadarlat–Eilers [12, 13]. It states that if a Cuntz pair $(\varphi,\psi)$ represents the zero element in $KK(A,B)$, then $\varphi$ and $\psi$ are homotopic via conjugating one of the maps with a unitary path in $\mathbf{1}+B\otimes\mathcal{K}$ after stabilizing both $\varphi$ and $\psi$ with a so-called absorbing representation. Note that an elegant treatment of the stable uniqueness theorem was recently included in [10]. Over the years, this theorem has become one of the central pieces of methodology in the structure and classification theory of nuclear $\mathrm{C}^{*}$-algebras; see for example [46, 26, 25, 16, 14, 22, 38, 23, 18, 24, 6]. More recently, I have generalized the stable uniqueness theorem to the setting of $\mathrm{C}^{*}$-dynamics together with Gabe in [21], which was one of the key conceptual ingredients behind our dynamical Kirchberg–Phillips theorem [20]. Further interesting applications of this theorem may be expected to come in the realm of $\mathrm{C}^{*}$-dynamics and their classification. Its counterpart for actions of unitary tensor categories followed soon afterwards [33] and may become instrumental to classify unitary tensor category actions on Kirchberg algebras.
When one considers the stable uniqueness theorem (either in its $\mathrm{C}^{*}$-algebraic or $\mathrm{C}^{*}$-dynamical form), the role of the absorbing representation is that of a necessary largeness criterion for the representations in question. The notion of an absorbing representation — recalled after subsection 1.1 for $\mathrm{C}^{*}$-dynamics — is at the same time inspired by the conclusion of Voiculescu’s fundamental theorem [47]. It can be interpreted as asserting that every unitizably full [^1] In this specific context, this can be interpreted as one of two more familiar statements. If the $\mathrm{C}^{*}$-algebra is non-unital, then the representation needs to be faithful and essential. If the $\mathrm{C}^{*}$-algebra is unital, then it also needs to have the property that the representation is supported on a subspace with infinite-dimensional orthogonal complement. representation of a separable $\mathrm{C}^{*}$-algebra on the Hilbert space $\ell^{2}(\mathbb{N})$ is absorbing. It is known that all elements of a given $KK$-group can be represented by a Cuntz pair of absorbing representations. When one is faced with applications to the classification of nuclear $\mathrm{C}^{*}$-algebras, such an absorbing behavior can be detected via Elliott–Kucerovsky’s “purely largeness” criterion [15] and its known variants; see also [3, Subsection 4.2]. For the purpose of this article, one of the primary applications to have in mind in classification is the approach of Carrión et al [6], where part of the technical work is about getting closer to a setup where one can compare a pair of naturally occurring absorbing representations via $KK$-theory. The stable uniqueness theorem allows one to obtain a strong form of equivalence for the pair of representations, but in general only after first stabilizing with an auxiliary representation. The process to “destabilize”, so to speak, may require a substantial amount of extra work involving the specific assumptions related to the application. Schafhauser demonstrated in [38] that this issue can be circumvented under a UHF-stability assumption. Building on this idea, it is shown in [6, Theorem 5.15] that one can omit the stabilization with the auxiliary representation at the expense of adding a tensor copy of the Jiang–Su algebra $\mathcal{Z}$ to the codomain. Since said article concerns the classification of $\mathrm{C}^{*}$-algebras that are ultimately Jiang–Su stable, this extra tensor copy of $\mathcal{Z}$ can be removed in a later phase of the application with some extra steps. However, it is a natural question to what extent it is really necessary to stabilize with an auxiliary representation when one compares a pair of representations that are absorbing to begin with. This was asked as [6, Question 5.17] and [39, Problem LXII]. One would refer to the desired statement below as a “$KK$-uniqueness theorem”.
Question A.
@@
Suppose that $A$ and $B$ are separable $\mathrm{C}^{*}$-algebras. Let $\varphi,\psi:A\to\mathcal{M}(B\otimes\mathcal{K})$ be two absorbing representations forming a Cuntz pair. If $\varphi,\psi]=0$ in $KK(A,B)$, is it true that $\varphi$ and $\psi$ are strongly asymptotically unitarily equivalent?
The discussions in the aforementioned references point towards the main technical obstacle being the $K_{1}$-injectivity of the auxiliary $\mathrm{C}^{*}$-algebra
$\mathcal{Q}(B\otimes\mathcal{K})\cap\bar{\varphi}(A)^{\prime},$
which had been a problem considered beforehand in other related works [37, 19, 35, 36] (here $\mathcal{Q}$” denotes the corona algebra and $\bar{\varphi}$ is the induced homomorphism). More recently, I demonstrated with Farah in [17, Theorem 5.5] that under the assumption of Jiang–Su stability on $B$, this auxiliary object is separably Jiang–Su stable, which implies $K_{1}$-injectivity (by [29] or [6, Theorem 4.8]) and thus the $KK$-uniqueness theorem is indeed true when the codomain algebra $B$ is Jiang–Su stable.
The first and main goal of this article is to answer Introduction in the affirmative, by way of showing that the technical obstacle related to $K_{1}$-injectivity has a positive solution for arbitrary pairs of separable $\mathrm{C}^{*}$-algebras. We mention briefly that as per the discussion around [6, Question 5.17], Introduction has a matching version involving a similar statement for $KL$-theory, where the desired conclusion is instead a version of approximate unitary equivalence of the two involved presentations. While this kind of $KL$-uniqueness is also very useful in applications, it is implicit from that discussion that it can be deduced as the consequence of a general answer to Introduction. This is the reason why $KL$-theory plays no explicit role in the rest of this article. To describe the approach to answer Introduction, it is more useful to consider the $\mathrm{C}^{*}$-algebra $\mathscr{D}_{\varphi}\subseteq\mathcal{M}(B\otimes\mathcal{K})$ of all multipliers whose induced class belongs to $\mathcal{Q}(B\otimes\mathcal{K})\cap\bar{\varphi}(A)^{\prime}$; see section 2. The main result of the second section is the following consequence of section 2 and section 2.
@@
Let $A$ and $B$ be separable $\mathrm{C}^{*}$-algebras. If $\varphi:A\to\mathcal{M}(B\otimes\mathcal{K})$ is (unitally) absorbing, then $\mathscr{D}_{\varphi}$ and all its quotients are $K_{1}$-injective.
Combined with the detailed discussions appearing after [6, Question 5.17] and [39, Problem LXII], this implies a positive answer to Introduction. The added layers of generality within section 2 include other recent results of this type like [35, Theorems 2.5, 2.9] and [36, Theorem 3.28] as special cases.
Before we move on to the remaining content of the article, I shall say a few words about the approach to prove Introduction. The core argument behind the proof draws heavy inspiration from the proof of Cuntz–Higson’s theorem [11] that the unitary group of every stable multiplier algebra $\mathcal{M}(B\otimes\mathcal{K})$ is connected in the norm topology. If one analyzes said proof carefully, one can interpret it as a two-step approach: The first (and relatively easy) step is to observe that the $K$-groups of $\mathcal{M}(B\otimes\mathcal{K})$ vanish, which is a consequence of the fact that it possesses an infinite repeat endomorphism, which automatically trivializes the identity map in $K$-theory. The second (and more difficult) step is to show that $\mathcal{M}(B\otimes\mathcal{K})$ is $K_{1}$-injective. Since it is properly infinite as a unital $\mathrm{C}^{*}$-algebra, there are known criteria to verify $K_{1}$-injectivity (these are nicely summarized in [2]), one of which was the original criterion verified by Cuntz–Higson. In essence, the proof of Introduction (say for the $K_{1}$-injectivity of the $\mathrm{C}^{*}$-algebra $\mathscr{D}_{\varphi}$ itself) can be seen to mimic Cuntz–Higson’s argument, with the added point that we need to keep track of the relative commutant condition appearing in the definition of $\mathscr{D}_{\varphi}$. Fortunately, this works in remarkable generality as long as $\varphi$ is sufficiently similar to an infinite repeat representation, which is automatically the case when $\varphi$ is absorbing. With this requisite in hand, the main difference between Cuntz–Higson’s original argument and the one appearing here is that instead of utilizing arbitrary increasing approximate units in $B\otimes\mathcal{K}$, we utilize approximate units that are quasicentral relative to $\varphi(A)$. Since the proof involves no further assumptions, the conclusion about $K_{1}$-injectivity can be drawn in a much more flexible situation (described in section 2), which is applicable to other auxiliary $\mathrm{C}^{*}$-algebras that play an analogous role for other types of $KK$-theory such as nuclear, ideal-related, or equivariant $KK$-theory. The argument does however very much depend on the fact that the $\mathrm{C}^{*}$-algebra in question comes from a stable multiplier algebra. Hence I would like to mention, for those readers wondering, that the proof behind Introduction likely sheds no light on the general $K_{1}$-injectivity problem for properly infinite $\mathrm{C}^{*}$-algebras; see [2, Question 2.9] and [39, Problem LXI].
@@
As an outlook for future applications, Hua–White’s recent article [28] serves as an excellent template. Therein, the authors prove a special case of the $KK$-uniqueness theorem, tailored to their intended application to classify certain nuclear embeddings of exact $\mathrm{C}^{*}$-algebras into ultrapowers of matrix algebras. Their setup did not match the assumptions required for the classification machinery in [6], but their work introduces various novel ideas that demonstrate how the scope of such classification results can be extended if one has access to the general $KK$-uniqueness theorem. I thus hope that the results of this article will be used in future works to obtain a more general classification of nuclear embeddings of exact $\mathrm{C}^{*}$-algebras into ultraproducts of $\mathrm{C}^{*}$-algebras that need not be Jiang–Su stable.

### v2

@@ Abstract.
Answering a question of Carrión et al in their recent landmark paper on $\mathrm{C}^{*}$-algebra classification, we prove a general uniqueness theorem for $KK$-theory. Given arbitrary separable $\mathrm{C}^{*}$-algebras $A$ and $B$ and a Cuntz pair consisting of two absorbing representations $\varphi,\psi:A\to\mathcal{M}(B\otimes\mathcal{K})$, the induced element of $KK(A,B)$ vanishes if and only if $\varphi$ and $\psi$ are strongly asymptotically unitarily equivalent. This improves upon the Lin–Dadarlat–Eilers stable uniqueness theorem. The conclusion is deduced by showing the $K_{1}$-injectivity of an auxiliary $\mathrm{C}^{*}$-algebra associated to the $\mathrm{C}^{*}$-pair $(A,B)$, which is sometimes called the Paschke dual algebra in the literature. Most of the article is concerned with the treatment of an umbrella theorem, which yields such a uniqueness theorem for other variants of $KK$-theory. This encompasses nuclear $KK$-theory, ideal-related $KK$-theory, equivariant $KK$-theory, or any combinations thereof.
@@ Introduction
Kasparov theory, also known as $KK$-theory, is a very powerful invariant for $\mathrm{C}^{*}$-algebras. It is a bivariant version of topological $K$-theory, which includes ordinary $K$-theory and $K$-homology, alongside Brown–Douglas–Fillmore theory [5], as special cases. Invented by Kasparov [29, 30], it can be viewed as either a bifunctor from (separable) C*-algebras to abelian groups, or as a way to equip the class of (separable) $\mathrm{C}^{*}$-algebras with a more flexible notion of homomorphisms. Given a second-countable locally compact group $G$, Kasparov’s construction also yields such a bifunctor on the category of all (separable) $G$-$\mathrm{C}^{*}$-algebras, which provided impactful applications on classical problems such as the Novikov conjecture. The original approach by Kasparov, commonly referred to as the Kasparov picture, is still the most common way of treating this subject, but there are other equivalent appproaches that are geared towards different types of applications.
In this article we are mainly interested in the so-called Cuntz picture [8] of $KK$-theory. In Cuntz’ framework, elements in Kasparov’s group $KK(A,B)$ for separable $\mathrm{C}^{*}$-algebras $A$ and $B$ can be described as homotopy classes of so-called Cuntz pairs, i.e., pairs of *-homomorphisms $\varphi,\psi:A\to\mathcal{M}(B\otimes\mathcal{K})$ such that $\varphi(a)-\psi(a)\in B\otimes\mathcal{K}$ for all $a\in A$. The vanishing of such a homotopy class can a priori be unraveled, roughly speaking, in terms of the existence of a particular kind of homotopy between the two representations $\varphi$ and $\psi$. In practice, one often deals with pairs of representations that are in a sense very large, which may allow for additional control over the specific form of such a homotopy. For applications of $KK$-theory to classification theory, which serves as the main motivation for this article, the best conclusion one can hope for is that one can find a continuous path of unitaries in $(B\otimes\mathcal{K})^{\dagger}$ that asymptotically conjugates $\varphi$ to $\psi$ in the point-norm topology. The constraint of the unitaries being chosen in $(B\otimes\mathcal{K})^{\dagger}$ is essential, since finding such a unitary path in the multiplier algebra is neither good enough for most applications, nor is it even sufficient to guarantee the vanishing of the $KK$-class associated to the pair $(\varphi,\psi)$ to begin with. Answering [6, Question 5.17] and [38, Problem LXII], this article aims to confirm that this best possible outcome is always attainable as soon as a known necessary condition is satisfied. One would refer to a statement of the form below as a “$KK$-uniqueness theorem”.
Theorem A.
@@
Suppose that $A$ and $B$ are separable $\mathrm{C}^{*}$-algebras and $B$ is stable. Let $\varphi,\psi:A\to\mathcal{M}(B)$ be two absorbing representations forming a Cuntz pair. If $\varphi,\psi]=0$ in $KK(A,B)$, then $\varphi$ and $\psi$ are strongly asymptotically unitarily equivalent, i.e., there exists a unitary path $v:[0,\infty)\to\mathcal{U}(\mathbf{1}+B)$ with $v_{0}=\mathbf{1}$ such that $\psi(a)=\lim_{t\to\infty}v_{t}\varphi(a)v_{t}^{*}$ for all $a\in A$.
Prior to this, the seminal theorem concerning the analytical properties of $KK$-theory (in the Cuntz picture) has been the so-called stable uniqueness theorem proved independently by Lin [33] and Dadarlat–Eilers [12, 13]. The theorem states that if an arbitrary Cuntz pair $(\varphi,\psi)$ represents the zero element in $KK(A,B)$, then $\varphi$ and $\psi$ are asymptotically unitarily equivalent as above (without demanding $v_{0}=\mathbf{1}$) after stabilizing both $\varphi$ and $\psi$ with an absorbing representation. It has developed into one of the central pieces of methodology in the structure and classification theory of nuclear $\mathrm{C}^{*}$-algebras; see for example [45, 26, 25, 16, 14, 22, 37, 23, 18, 24, 6]. Note that an elegant treatment of the stable uniqueness theorem was recently included in [10].
In the context of both Introduction and the prior stable uniqueness theorem, the role of the absorbing assumption is that of a necessary largeness criterion for the representations in question. The notion of an absorbing representation — recalled after subsection 1.1 for $\mathrm{C}^{*}$-dynamics — is inspired by the conclusion of Voiculescu’s fundamental theorem [46]. Its assertion is equivalent to saying that every unitizably full [^1] In this specific context, this becomes one of two more familiar statements. If the $\mathrm{C}^{*}$-algebra is non-unital, then the representation needs to be faithful and essential. If the $\mathrm{C}^{*}$-algebra is unital, then it also needs to have the property that the representation is supported on a subspace with infinite-dimensional orthogonal complement. representation of a separable $\mathrm{C}^{*}$-algebra on the Hilbert space $\ell^{2}(\mathbb{N})$ is absorbing. Absorbing representations always exist between separable $\mathrm{C}^{*}$-algebras [43] and it is known that all elements of a given $KK$-group can be represented by a Cuntz pair of absorbing representations. When one is faced with applications to the classification of nuclear $\mathrm{C}^{*}$-algebras, one can detect such absorbing behavior via Elliott–Kucerovsky’s “pure largeness” criterion [15] and its known variants; see [3, Subsection 4.2]. For the purpose of this article, the primary application to have in mind in classification is the approach of Carrión et al [6], where part of the technical work is about getting closer to a setup where one can compare a pair of naturally occurring absorbing representations via $KK$-theory. The stable uniqueness theorem allows one to obtain a strong form of equivalence for the pair of representations, but only after first stabilizing with an auxiliary representation. The process to “destabilize”, so to speak, often required a substantial amount of extra work involving the specific assumptions related to the application. Schafhauser demonstrated in [37] how this issue can be circumvented under a UHF-stability assumption. Building on this idea, it was shown in [6, Theorem 5.15] that one can omit the stabilization with the auxiliary representation at the expense of adding a tensor copy of the Jiang–Su algebra $\mathcal{Z}$ to the codomain. Since said article concerns the classification of $\mathrm{C}^{*}$-algebras that are ultimately Jiang–Su stable, this extra tensor copy of $\mathcal{Z}$ can be removed in a later phase of the application with some extra steps. More recently, I strengthened said result with Farah in [17, Theorem 5.5] by verifying that Introduction holds under the extra assumption of Jiang–Su stability on $B$. However, it remained a natural question to what extent it is generally necessary to stabilize with an auxiliary representation when one compares a pair of representations that are absorbing to begin with, to which Introduction provides the answer.
If one broadens the scope of this investigation to equivariant $KK$-theory, there is an analog of Cuntz’ picture involving pairs of cocycle representations satisfying a similar condition as above; see also [40, Section 6]. Similarly to the nonequivariant case, the Cuntz–Thomsen picture [41] for equivariant $KK$-theory has turned out to be very useful for the classification of $\mathrm{C}^{*}$-dynamics. I have generalized the stable uniqueness theorem to the setting of $\mathrm{C}^{*}$-dynamics together with Gabe in [21], which was one of the key conceptual ingredients behind our dynamical Kirchberg–Phillips theorem [20]. Its counterpart for actions of unitary tensor categories followed soon afterwards [32] and may become instrumental to classify unitary tensor category actions on Kirchberg algebras.
The discussions in the aforementioned references [6, 38] identified the main technical obstacle towards the conclusion of Introduction as the $K_{1}$-injectivity of the auxiliary $\mathrm{C}^{*}$-algebra $\mathcal{Q}(B)\cap\bar{\varphi}(A)^{\prime}$, which had been a problem considered beforehand in other related works [36, 19, 34, 35] (here $\mathcal{Q}$” denotes the corona algebra and $\bar{\varphi}$ is the induced homomorphism). The first and main goal of this article is to prove Introduction, by way of showing that the technical obstacle related to $K_{1}$-injectivity has a positive solution for arbitrary pairs of separable $\mathrm{C}^{*}$-algebras. To describe the approach behind this, it is more useful to consider the $\mathrm{C}^{*}$-algebra $\mathscr{D}_{\varphi}\subseteq\mathcal{M}(B)$ of all multipliers whose induced class belongs to $\mathcal{Q}(B)\cap\bar{\varphi}(A)^{\prime}$; see section 2. The main result of the second section is the following consequence of section 2 and section 2. The added layers of generality within section 2 include other recent results of this type like [34, Theorems 2.5, 2.9] and [35, Theorem 3.28] as special cases.
@@
Let $A$ and $B$ be separable $\mathrm{C}^{*}$-algebras with $B$ stable. If $\varphi:A\to\mathcal{M}(B)$ is a (unitally) absorbing *-homomorphism, then $\mathscr{D}_{\varphi}$ and all its quotients are $K_{1}$-injective.
Before we move on to the remaining content of the article, I shall say a few words about the ideas leading to Introduction. The core argument behind the proof draws heavy inspiration from the proof of Cuntz–Higson’s theorem [11] that the unitary group of every stable multiplier algebra $\mathcal{M}(B)=\mathcal{M}(B\otimes\mathcal{K})$ (recal that $B$ is assumed stable) is connected in the norm topology. If one analyzes said proof carefully, one can interpret it as a two-step approach: The first (and relatively easy) step is to observe that the $K$-groups of $\mathcal{M}(B)$ vanish, which is a consequence of the fact that it possesses an infinite repeat endomorphism, which automatically trivializes the identity map in $K$-theory. The second (and more difficult) step is to show that $\mathcal{M}(B)$ is $K_{1}$-injective. Since it is properly infinite as a unital $\mathrm{C}^{*}$-algebra, there are known criteria to verify $K_{1}$-injectivity (these are nicely summarized in [2]), one of which was the original criterion verified by Cuntz–Higson. The proof of Introduction (say for the $K_{1}$-injectivity of the $\mathrm{C}^{*}$-algebra $\mathscr{D}_{\varphi}$ itself) can be seen to mimic Cuntz–Higson’s argument, with the added point that we need to keep track of the relative commutant condition appearing in the definition of $\mathscr{D}_{\varphi}$. Fortunately, this works in remarkable generality as long as $\varphi$ is sufficiently similar to an infinite repeat representation, which is automatically the case when $\varphi$ is absorbing. With this requisite in hand, the main difference between Cuntz–Higson’s original argument and the one appearing here is that instead of utilizing arbitrary increasing approximate units in $B$, we utilize approximate units that are quasicentral relative to $\varphi(A)$. Since the proof involves no further assumptions, the conclusion about $K_{1}$-injectivity can be drawn in a much more flexible situation (described in section 2), which is applicable to other auxiliary $\mathrm{C}^{*}$-algebras that play an analogous role for other types of $KK$-theory such as nuclear, ideal-related, or equivariant $KK$-theory. The argument does however very much depend on the fact that the $\mathrm{C}^{*}$-algebra in question comes from a stable multiplier algebra. Hence I would like to mention, for those readers wondering, that the proof behind Introduction likely sheds no light on the general $K_{1}$-injectivity problem for properly infinite $\mathrm{C}^{*}$-algebras; see [2, Question 2.9] and [38, Problem LXI].
@@
Upon applying this umbrella theorem to more familiar cases, we deduce Corollaries 4, 4 and 4, which are variants of Introduction for equivariant $KK$-theory as well as nuclear and/or ideal-related $KK$-theory.
@@
As an outlook for future applications of the main results, Hua–White’s recent article [28] serves as an excellent template. Therein, the authors prove a special case of the $KK$-uniqueness theorem, tailored to their intended application to classify certain nuclear embeddings of exact $\mathrm{C}^{*}$-algebras into ultrapowers of matrix algebras. Their setup did not match the assumptions required for the classification machinery in [6], but their work introduces a streamlined framework to extend the scope of such classification results if one has access to the general $KK$-uniqueness theorem. I thus expect that Introduction and its variants will be used in future works to obtain a more general classification of nuclear embeddings of exact $\mathrm{C}^{*}$-algebras into ultraproducts of $\mathrm{C}^{*}$-algebras that need not be Jiang–Su stable. This expectation appears to be supported by the discussion in [6, 1.3.6]. Similar consequences can be predicted within the classification theory of $\mathrm{C}^{*}$-dynamics.